### v1

# The Art of Imitation: Learning Long-Horizon Manipulation Tasks from Few Demonstrations

Abstract
Task Parametrized Gaussian Mixture Models (TP-GMM) are a sample-efficient method for learning object-centric robot manipulation tasks. However, there are several open challenges to applying TP-GMMs in the wild. In this work, we tackle three crucial challenges synergistically. First, end-effector velocities are non-Euclidean and thus hard to model using standard GMMs. We thus propose to factorize the robot’s end-effector velocity into its direction and magnitude, and model them using Riemannian GMMs. Second, we leverage the factorized velocities to segment and sequence skills from complex demonstration trajectories. Through the segmentation, we further align skill trajectories and hence leverage time as a powerful inductive bias. Third, we present a method to automatically detect relevant task parameters per skill from visual observations. Our approach enables learning complex manipulation tasks from just five demonstrations while using only RGB-D observations. Extensive experimental evaluations on RLBench demonstrate that our approach achieves state-of-the-art performance with 20-fold improved sample efficiency. Our policies generalize across different environments, object instances, and object positions, while the learned skills are reusable.
I Introduction
Learning robot manipulation tasks from a handful of demonstrations is a challenging problem [1, 2]. This problem is exacerbated when learning only from visual observations and aiming to generalize across task instances. Deep learning approaches can be parameterized flexibly and generalize well, but usually require hundreds of demonstrations or multiple stages [3]. Leveraging compact observation spaces, such as scene keypoints, alleviates the problem, but does not resolve it entirely [4]. In contrast, Gaussian Mixture Models (GMMs) [5] can be learned from only a handful of demonstrations. Various extensions to the GMM framework exist to enable generalization across task instances. For example, Task-Parameterized GMMs (TP-GMM) [6] generalize by modeling the set of demonstrations from the perspective of multiple relevant coordinate frames, called task-parameters.
Figure 1: TAPAS-GMM: Task Auto-Parameterized And Skill Segmented GMM learns task-parameterized manipulation policies from only a handful of complex task demonstrations. First, we segment the full task demonstrations into the involved skills. For each segment, we then automatically select the relevant task parameters and learn a Riemannian Task-Parameterized Hidden Markov Model (TP-HMM). The skill models can be cascaded and reused flexibly. To enable modeling of the robot’s end-effector velocity, we further leverage a novel action factorization and Riemannian geometry.
Current methods based on TP-GMMs exhibit a set of serious limitations. First, velocity information is often important for effectively modeling robotic tasks. However, velocities are best expressed by their direction and magnitude, and are poorly represented using Euclidean approaches. As GMMs typically assume Euclidean data, velocity information is difficult to model using TP-GMMs [7]. Second, GMMs suffer from the curse of dimensionality, i.e. they scale poorly to both high-dimensional data and long trajectories. Both are crucial drawbacks in a robotics context. On the one hand, the dimensionality of the TP-GMM scales linearly with the number of task parameters. On the other hand, even seemingly simple tasks such as pick and place can be too long and complex to fit. Third, tasks often consist of multiple skills such as pick up the bottle. These skills might not be temporally aligned across demonstrations, rendering the data distribution even more complex. Fourth, when learning from visual observations, generating a set of candidate task parameters is an open problem [8, 9]. In practice, object tracking systems are used [10], but this hardware dependency severely limits the model’s applicability. Finally, the task-parameter selection is crucial for the resulting model quality. However, existing approaches for selecting among a set of candidates also suffer from the curse of dimensionality [11].
We propose Task Auto-Parameterized And Skill Segmented GMM TAPAS-GMM, illustrated in Fig. 1. Our method addresses this set of challenges synergistically. First, we propose to refactorize the robot’s velocity into direction and magnitude to better capture the underlying data distribution. We then model both using Riemannian GMMs [10] which are not constrained by the Euclidean assumptions. Second, we segment the task demonstrations into skills using the velocity information. The segmentation both aligns the trajectories temporally and enables learning skill models with shorter trajectories and lower data dimensionality. Additionally, the temporal alignment enables us to use time as a powerful inductive bias for the initialization of our model. If the trajectories are aligned, then simple binning of the trajectories over the time dimension yields an excellent initialization. Third, we use a pretrained visual encoder network [12] to generate object keypoints that generalize across object instances and task environments [4]. This way, we generate a set of candidate task parameters from RGB-D observations. Finally, for selecting the task-relevant parameters from these candidates, we offer an approach that scales well to large numbers of candidates. Our proposed solutions have synergistic effects. For example, combining our parameter selection and skill segmentation further reduces the data dimensionality, as a given task parameter is often only relevant for some skill and not the whole long-horizon task.
We evaluate our method on a diverse set of manipulation tasks, both in RLBench [13] and on a real robot, including articulated objects, as well as high-precision, and long-horizon tasks. We experimentally disentangle policy learning and representation learning, as well as ablating the major components of our approach. TAPAS-GMM learns long-horizon manipulation policies from as little as five demonstrations. We show generalization across object positions, object instances, visual clutter, and task environments. Finally, we demonstrate how the learned skills can be recombined in novel ways.
In summary, our main contributions are
  1. 1.
    a novel action factorization to effectively model end-effector velocities using Task-Parameterized Gaussian Mixture Models.
  2. 2.
    a new velocity-based approach to segmenting complex demonstration trajectories into skills.
  3. 3.
    an approach that efficiently extracts and selects task parameters per skill from visual observations.
  4. 4.
    rigorous evaluation of our approach’s ability to overcome the problems outlined above, and its efficacy for policy learning, both in simulation and on a real robot.
  5. 5.
    We make the code and models publicly available at http://tapas-gmm.cs.uni-freiburg.de.
The supplementary material is appended at the end of this paper.
II Related Work
Parameterized Motions: GMMs [5] learn robotic manipulation tasks from only a handful of demonstrations. Popular alternatives are Dynamic Motion Primitives (DMP) [14] and Probabilistic Motion Primitives (ProMP) [15]. However, all three methods require additional techniques to adapt to novel task instances. Calinon et al. [6] identify three classes of adaptation methods: model database, parametric, and multi-stream, to which we amend two-stage and point-based approaches. Model database approaches [16] learn one model per demonstration, that they retrieve and combine based on the associated parameters. In contrast, parametric models [17] learn a single, parameterized model. Two-stage approaches first learn a model encoding a primitive motion, that they adapt in a second stage, e.g. using reinforcement learning [18] or Wasserstein gradients [19]. However, this requires a second online learning phase. Whereas, multi-stream approaches [20] model the same set of demonstrations from the perspective of multiple coordinate frames, called task-parameters. They generalize via the transformation and combination of these local models. Task-parameterized Gaussian Mixture Models (TP-GMM) [6, 10, 21] are multi-stream methods, leveraging a single expectation-maximization procedure. Finally, point-based approaches [9, 22] warp the underlying model to explicit endpoints and waypoints.
TP-GMMs estimate the relevance of a task parameter for a given observation from the variance observed in the task demonstrations. Consequently, they generalize poorly when given fewer than a handful of demonstrations [9]. In contrast, directly editing waypoints [9, 22] is feasible for just a single demonstration. However, task parameters can be extracted from visual observations [23], while defining waypoints may require significant engineering effort. Moreover, by using multiple demonstrations, TP-GMMs approximate the full trajectory distribution suited to solve a given task, instead of assuming the optimality of a single demonstration. Most importantly, TP-GMMs can be extended to model rotations [10], which are required to effectively solve many robotic manipulation tasks. Whereas, both Elastic-DS [9] and KMP [22] are only defined for Euclidean spaces.
Parameter Selection: Task-parameters are usually “assumed to be known or given during demonstrations” [9]. This requires “the experimenter to provide an initial set of [potentially relevant candidate] frames” [8]. Given a set of candidate frames, their respective relevance for a given demonstration can be estimated [24]. However, current approaches fit a joint model to all candidate frames [24]. Hence, they suffer from the curse of dimensionality, which our method avoids. Furthermore, automatically generating an initial set of candidate frames is an open problem [8, 9]. In practice, infrared motion tracking devices are manually attached to relevant objects [10], limiting the practical applicability of these methods. In contrast, we propose to automatically identify candidate task parameters from visual observations and to select those relevant to the task at hand. This enables instance-level generalization across objects and applications in novel and changing environments.
Sample-Efficient Long-Horizon Manipulation: GMMs are typically used to model full trajectories [18], with some work focusing on solving long-horizon tasks by sequencing learned GMMs in a bottom-up manner [21]. In contrast, we focus on directly learning long-horizon tasks by segmenting and sequencing skills top-down from complex task demonstrations. Some deep learning based approaches [25, 26] simplify long-horizon problems, by only predicting a set of key poses and using motion planning in between. However, by throwing away the trajectory information between key poses, they cannot produce constraint motions such as opening a cabinet. Moreover, they can only predict key poses in the demonstrated order, whereas our approach further produces a set of reusable and re-combinable skills.
III Background
In this section, we summarize foundational work on Gaussian Mixture Models for robotic manipulation, including Hidden Markov Models (HMM), task-parameterization, and Riemannian GMMs. We provide additional details including full equations in Sec. S.1 of the supplementary material.
[…]
III-A Gaussian Mixture Model
A Gaussian Mixture Model (GMM) with $K$ Gaussian components is defined as $\mathcal{G}=\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ denote the $k$-th component’s mean and covariance matrix and $\pi_{k}$ denotes its prior probability. In the context of Learning from Demonstrations, GMMs are employed to estimate the joint distribution of a disjoint set of input variables $\mathcal{I}$ and output variables $\mathcal{O}$, from a set of task demonstrations $\mathcal{D}$ as
$p_{\mathcal{G}}(\mathcal{D})=p(\mathcal{I},\mathcal{O}\mid\mathcal{G})=\sum_{k=1}^{K}\pi_{k}\cdot\mathcal{N}(\mathcal{I},\mathcal{O}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}).$ (1)
We define $\mathcal{D}=\{\{\boldsymbol{z}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$, where $\boldsymbol{z}\in\mathcal{I}\cup\mathcal{O}$, while $t$ indexes time and $n$ the demonstration. We use Expectation Maximization (EM) to maximize the probability of $\mathcal{D}$ under $\mathcal{G}$. For inference, we employ Gaussian Mixture Regression (GMR) [27] to estimate the conditional density
$p_{\mathcal{G}}(\boldsymbol{o}\mid\boldsymbol{i})=\sum_{k=1}^{K}\pi_{k}^{\boldsymbol{i}}\cdot\mathcal{N}(\boldsymbol{o}\mid\boldsymbol{\mu}_{k}^{o\mid i},\boldsymbol{\Sigma}_{k}^{o\mid i}).$ (2)
Next, we discuss the choice of input and output variables.
III-B Time-Driven vs. State-Driven Models
There are at least two ways to model robotic GMM policies: time-driven and state-driven. Let $\boldsymbol{\xi}$ denote the robot’s end-effector pose. Time-driven GMMs model the robot’s state conditional on time: $p(\boldsymbol{\xi}\mid t)$, i.e. =$\mathcal{I}=\{t\}$ and $\mathcal{O}=\{\boldsymbol{\xi}\}$. As they do not model the robot’s velocity, we post-process the predictions to ensure compliance with the robot’s velocity limits, e.g. using TOPPRA [28].
State-driven GMMs model the robot’s velocity given its current state: $p(\dot{\boldsymbol{\xi}}\mid\boldsymbol{\xi})$, i.e. =$\mathcal{I}=\{\boldsymbol{\xi}\}$ and $\mathcal{O}=\{\dot{\boldsymbol{\xi}}\}$, which also captures the robot’s velocity limits. However, velocities are harder to model because they are non-Euclidean. Specifically, velocity trajectories are usually less smooth and have more multimodalities than state trajectories. Moreover, velocity errors can aggregate over time. Additionally, deployment on a robot can require time-independence of the policy, e.g. to effectively deal with disturbances. We evaluate these aspects experimentally in Sec. V-B.
III-C Hidden Markov Model
The GMM’s posterior $p_{\mathcal{G}}(\boldsymbol{o}\mid\boldsymbol{i})$ in Eq. (2) only depends on the priors $\pi_{k}^{i}$ and the conditioned Gaussians, not on any previous observations. The Hidden Markov Model (HMM) [29] augments the GMM by associating each Gaussian with a state $k$ and modeling the transition probability $a_{kl}$ from state $k$ to state $l$. I.e. =$\mathcal{H}=\{\pi_{k},\{a_{kl}\}_{l=1}^{K},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$. This helps the model to differentiate between components that are close together in the input space, yet model different parts of a task or movement.. For a sequence of observations $\{\boldsymbol{i}_{t}\}_{t=0}^{T}$, the computation of the prior thus changes to
$\pi_{k}^{i_{t}}=\frac{\left(\sum_{l=1}^{K}\pi_{l}^{i_{t-1}}a_{lk}\right)\mathcal{N}(\boldsymbol{i}_{t}\mid\boldsymbol{\mu}_{k}^{\mathcal{I}},\boldsymbol{\Sigma}_{k}^{\mathcal{I}})}{\sum_{m=1}^{K}\left[\left(\sum_{l=1}^{K}\pi_{l}^{i_{t-1}}a_{lm}\right)\mathcal{N}(\boldsymbol{i}_{t}\mid\boldsymbol{\mu}_{m}^{\mathcal{I}},\boldsymbol{\Sigma}_{m}^{\mathcal{I}})\right]}.$ (3)
For clarity, we continue referring to our models as GMMs, even though all of them use the HMM extension.
Figure 3: Velocity trajectories from the first skill in StackWine. Left: The velocities are difficult to cluster in Euclidean space. Middle: $\mathcal{S}^{2}$ models the movement direction. Right: The associated action magnitudes.
III-D Task-Parameterization
Task-Parameterized GMMs [6] model the movement from the perspective of multiple coordinate frames, called task-parameters, in one joint GMM [10]. These frames are attached to task-relevant objects, but may also include the robot’s frame and a world frame. Let $\boldsymbol{x}(f)$ denote the end-effector position in the frame of task-parameter $f$. We discuss the extension to full poses $\boldsymbol{\xi}(f)$ in Sec. III-E. For the time-driven model and $F$ task-paramaters, we model =$\mathcal{I}=\{t\}\text{ and }\mathcal{O}=\{\boldsymbol{x}(f)\}_{f=1}^{F}$. In contrast, for the state-driven model, we define =$\mathcal{I}=\{\boldsymbol{x}(f)\}_{f=1}^{F}\text{ and }\mathcal{O}=\{\dot{\boldsymbol{x}}(f)\}_{f=1}^{F}$. Akin to the vanilla GMM, we fit all parameters jointly using EM. Next, we compute the per-frame marginal model for each frame $1\leq f\leq F$. For example, in the state-driven case we estimate
$p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G})=p(\boldsymbol{x}(f),\dot{\boldsymbol{x}}(f)\mid\mathcal{G}).$ (4)
To adapt to a given set of task parameters $\Theta=\{\boldsymbol{A}_{f},\boldsymbol{b}_{f}\}_{f=1}^{F}$, we transform the per-frame marginals to the world frame. $\boldsymbol{b}_{f}$ denotes the origin of frame $f$ and $\boldsymbol{A}_{f}$ its rotation matrix. Let
$p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G})=\left\{\pi_{k}^{f},\boldsymbol{\mu}_{k}^{f},\boldsymbol{\Sigma}_{k}^{f}\right\}_{k=1}^{K}$ (5)
be a marginal. Then its transformed marginal is given by
$\begin{split}p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G},\Theta)&=\left\{\pi_{k}^{f},\hat{\boldsymbol{\mu}}^{f}_{k},\hat{\boldsymbol{\Sigma}}^{f}_{k}\right\}_{k=1}^{K}\\ \hat{\boldsymbol{\mu}}^{f}_{k}&=\boldsymbol{A}_{f}\boldsymbol{\mu}_{k}^{f}+\boldsymbol{b}_{f}\\ \hat{\boldsymbol{\Sigma}}^{f}_{k}&=\boldsymbol{A}_{f}\boldsymbol{\Sigma}_{k}^{f}\boldsymbol{A}_{f}^{T}.\end{split}$ (6)
Finally, we compose the joint model as
$p_{\mathcal{G}}(\mathcal{I},\mathcal{O}\mid\Theta)=\prod_{f=1}^{F}p_{f}(\mathcal{I},\mathcal{O}\mid\mathcal{G},\Theta).$ (7)
We then perform GMR on the joint model as defined in Eq. (2). For an excellent, more detailed overview of TP-GMMs, please refer to [8].
III-E Riemannian Manifolds
We model the robot’s end-effector pose $\boldsymbol{\xi}=\begin{bmatrix}\boldsymbol{x}&\boldsymbol{\theta}\end{bmatrix}^{T}$ on the Riemmanian Manifold $\mathcal{M}_{\text{pose}}=\mathbb{R}^{3}\times\mathcal{S}^{3}$. Whereas, GMMs typically assume Euclidean data, the naive Euclidean mean of a set of quaternions is generally not a valid quaternion. Instead of projecting the naive solution back onto the manifold, we solve the problem rigorously using Riemannian geometry [10]. Each Riemannian manifold $\mathcal{M}$ is differentiable, i.e., at each point $\mathbf{x}$, it has a tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}$, that is a real vector space. The logarithmic map $\text{Log}_{\mathbf{x}}:\mathcal{M}\to\mathcal{T}_{\mathbf{x}}\mathcal{M}$ straightens the manifold’s curvature while preserving geodesic distance. The inverse is the exponential map $\text{Exp}_{\mathbf{x}}:\mathcal{T}_{\mathbf{x}}\mathcal{M}\to\mathcal{M}$. Both depend on the base $\mathbf{x}$, at which they are applied. Following Zeestraten [10], we formulate all previously Euclidean operations in EM and GMR as non-linear Maximum Likelihood Estimation problems. For example, to compute the center of mass $\boldsymbol{\mu}$ of a Riemannian Gaussian, we initialize the procedure with some estimate $\hat{\boldsymbol{\mu}}$ and solve the MLE using the following iterative process.
  1. 1.
    Tangent project the data using $\text{Log}_{\hat{\boldsymbol{\mu}}}$.
  2. 2.
    Update $\Delta\in\mathcal{T}_{\hat{\boldsymbol{\mu}}}\mathcal{M}$ using a Gauss-Newton step.
  3. 3.
    Back-project the update $\hat{\boldsymbol{\mu}}\leftarrow\text{Exp}_{\hat{\boldsymbol{\mu}}}(\Delta)$.
In practice, only a handful of iterations suffice for $\hat{\boldsymbol{\mu}}$ to converge. Fig. 2 illustrates the procedure. All other operations such as conditioning can be solved in a similar manner [10]. Finally, we use the concept of parallel transport to implement the affine frame transformations in Eq. (6) for the full pose. We provide additional details in Sec. S.2 of the supplementary material. Zeestraten provides an excellent introduction into TP-GMMs on Riemannian Manifolds [10]. Note, that we found parallel transport to sometimes introduce unwanted correlations into the model’s covariance, leading to faulty predictions. We discuss the problem and our proposed solution in detail in Sec. S.3 of the supplementary material.
Figure 4: Left: GMM on unaligned demos. Right: GMM on aligned demos. Dotted lines indicate segment borders.
IV Technical Approach
We aim to learn robot manipulation policies that generalize to unseen task instances from just five demonstrations. To generalize across object instances and be independent of object-tracking hardware, the task parameters should be extracted from RGB-D observations. For fast and easy data collection, complex tasks should automatically be segmented into their atomic skills. The method should not require one collection phase per atomic skill and should be able to construct novel sequences that were not demonstrated.
We present TAPAS-GMM: Task-Parameterized and Skill Segmented GMM. Our approach is based on Riemannian Task-Parameterized Hidden Markov Models. To effectively model velocity data, we propose a novel factorization. We then segment skills from complex demonstrations to temporally align the demonstrations and to learn skill models that we can sequence and reuse later. Finally, we propose an approach to automatic task parameterization and combine these techniques in a synergistic manner. Fig. 5 summarizes our approach. We now introduce the components in turn.
Figure 5: Overview of our approach. Learning: First, we segment a set of complex task demonstrations with unaligned skills. Next, we generate a set of candidate task parameters from visual observations and select the relevant parameters for each segment. Finally, we fit one Task-Parameterized Hidden Markov Model (TP-HMM) per segment. Inference: To make a prediction for new visual observations, we again extract the set of task parameters and select the task parameters determined during the learning phase. We then cascade the segment TP-HMMs.
Figure 6: Left: The RLBench tasks OpenDrawer, PushButtons, StackWine, SlideBlock, SweepToDustpan, and PlaceCups. Right: The real-world tasks PickAndPlace, PourDrink, MakeCoffee, SweepBlocks, OpenCabinet, StoreInDrawer, and BakeBread.
IV-A Action Factorization
In Sec. III-B, we discussed that solving manipulation tasks can require modeling the end-effector’s velocity $\dot{\boldsymbol{\xi}}$ and not only its pose $\boldsymbol{\xi}$, e.g. if time-independence is required. Fig. 3 illustrates that naively modeling $\dot{\boldsymbol{\xi}}$ on the manifold $\mathcal{M}_{\text{pose}}=\mathbb{R}^{3}\times\mathcal{S}^{3}$ (introduced in Sec. III-E) is difficult because velocities are not Euclidean. Instead, they are vector quantities defined by a direction and magnitude. We therefore propose to factorize the end-effector velocity $\dot{\boldsymbol{\xi}}$ as
$\dot{\boldsymbol{\xi}}=\begin{bmatrix}\hat{\dot{\boldsymbol{x}}}&\hat{\dot{\boldsymbol{\theta}}}&\left\lVert\dot{\boldsymbol{x}}\right\rVert&\left\lVert\dot{\boldsymbol{\theta}}\right\rVert\end{bmatrix}^{T}.$ (8)
The translation direction $\hat{\dot{\boldsymbol{x}}}$ is the normalized translation vector, whereas $\hat{\dot{\boldsymbol{\theta}}}$ is the rotation axis. Both lie on $\mathcal{S}^{2}$. The translation magnitude $\left\lVert\dot{\boldsymbol{x}}\right\rVert\in\mathbb{R}$ is the norm of the translation vector and $\left\lVert\dot{\boldsymbol{\theta}}\right\rVert\in\mathcal{S}^{1}$ is the rotation angle. The combined manifold is $\mathcal{M}_{\text{vel}}=\mathcal{S}^{2}\times\mathcal{S}^{2}\times\mathbb{R}\times\mathcal{S}^{1}$. For a set $F$ of task parameters, the full manifold of our state-driven TP-GMM is given by
$\mathcal{M}=(\mathcal{M}_{\text{pose}}\times\mathcal{S}^{2}\times\mathcal{S}^{2})^{F}\times\mathbb{R}\times\mathcal{S}^{1}.$ (9)
Our insight is three-fold. First, in the task-parameterized setting, only the direction of the velocities varies between different fixed coordinate frames, not their magnitude. Thus, we reduce the dimensionality of the problem by modeling the magnitudes globally. Second, because velocities are not Euclidean, their distributions are easier to estimate when factorized. To model the factorized velocities, we leverage the Riemannian toolbox introduced in Sec. III-E. Third, we leverage the isolated velocities to automatically segment complex task trajectories into the involved skills. We found this to drastically improve the quality and transferability of the learned model and discuss it in detail in Sec. IV-C. Fig. 3 illustrates how the velocities are easier to model in the factorized representation. We provide additional details on how to implement the frame transformations for $\mathcal{S}^{2}$ in Sec. S.2 of the supplementary material.
IV-B Gripper Action
Effectively solving manipulation tasks also requires modeling the gripper action. To this end, we add another global action dimension $\mathbb{R}$, modeling the gripper width, to Eq. (9).
IV-C Skill Segmentation
Robot manipulation tasks often involve a set of skills. The pick and place task in Fig. 5 consists of four skills. 1. Align the gripper with the first object. 2. Grasp it. 3. Align the first object with the second object. 4. Place the first object. All skills have distinct dynamics, so it is difficult to model them jointly in a single state-driven GMM. Furthermore, the skills have different relative durations across task instances, leading to poor temporal alignment across demonstrations. Fig. 4 illustrates that as a consequence they are difficult to model with a single time-driven GMM.
Therefore, we propose to automatically segment the full demonstration trajectories into a sequence of skill demonstrations and to learn one model per skill. This allows us to temporally align the skills and to learn local dynamics, specific to the skills. The skill model can be time-driven or state-driven. Furthermore, it allows us to reuse the learned skills by sequencing them in novel ways. Lastly, in Sec. IV-E and Sec. IV-F we discuss its synergistic effects with our initialization strategy and our automatic task-parameterization, respectively. To segment the trajectories, we threshold the action magnitudes introduced in Sec. IV-A. Observations with near-zero action magnitude yield a set of segmentation point candidates that we then cluster across the set of demonstrations using density-based clustering. Afterward, we fit one TP-HMM per skill.
IV-D Skill Sequencing and Skill Reuse
We execute a sequence of skills by cascading the associated HMMs. To cascade two HMMs $\mathcal{H}_{1},\mathcal{H}_{2}$, Rozo et al. astutely propose to calculate the transition probability $a_{ij}$ between state $i\in\mathcal{H}_{1}$ and state $j\in\mathcal{H}_{2}$ using the KL-divergence between the associated Gaussians[21] as
$a_{ij}\propto\exp\left(-D_{KL}(\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\parallel\mathcal{N}(\boldsymbol{\mu}_{j},\boldsymbol{\Sigma}_{j}))\right).$ (10)
In case $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ contain different task parameters, we marginalize the Gaussians to their set of common task parameters first. Afterward, we normalize the outgoing probabilities of all states. Computing the KL-divergence can be difficult for Riemannian Gaussians, so we estimate it efficiently using a Monte-Carlo sampling scheme [30]. We employ this procedure iteratively to longer sequences of HMMs. Note that this skill sequencing procedure does not require the skills to be sequenced in the demonstrated order. Instead, it allows us to solve new tasks by reordering and sequencing learned skills. Note further that time-based GMM skills can easily be reversed, either by manipulating the mean and covariance matrices or by reversing the time signal. In combination, this allows us to create a rich set of new task policies, which we demonstrate in Sec. V-D.
TABLE I: Success rates of the learned policies in simulated tasks. Results for 3D Diffuser Actor were reported by the authors.
Demonstrations Method Task
OpenDrawer PushButtons StackWine SlideBlock SweepToDustpan PlaceCups
100 3D Diffuser Actor [31] 0.90 0.98 0.93 0.98 0.84 0.24
Diffusion Policy [32] 0.96 0.00 0.42 0.31 0.32 0.00
LSTM [33] 0.00 0.06 0.00 0.00 0.00 0.00
5 Diffusion Policy 0.16 0.00 0.00 0.05 0.40 0.00
LSTM 0.00 0.00 0.00 0.00 0.00 0.00
TAPAS-GMM (Ours) 0.90 1.00 1.00 0.95 1.00 0.75
TABLE II: Policy success without task-parameter selection or skill segmentation. TP-GMM uses neither. The other models ablate one factor each.
OpenDrawer SlideBlock PlaceCups
TAPAS-GMM (Ours) 0.90 0.95 0.75
└─ w/o segmentation 0.06 0.68 0.00
└─ w/o TP selection 0.62 0.66 0.03
└─ Global TP selection 0.62 0.66 0.10
TP-GMM 0.05 0.64 0.00
IV-E Time-Based Initialization
EM is not guaranteed to converge to the global optimum, but only to a local optimum [10]. Consequently, its initialization is of critical importance. $K$-means clustering is popular but does not guarantee a global optimum either [10]. We efficiently segment the trajectories along the time dimension using $K$-bins. Fig. 4 illustrates that complex task demonstrations are usually not aligned temporally. However, our skill segmentation proposed in Sec. IV-C aligns them, yielding a good initialization for EM. The combination of skill segmentation and time-based initialization enables the EM procedure to arrive at good local optima, while also achieving significantly faster convergence. Next, we discuss how we leverage it to efficiently select task parameters, too.
IV-F Task-Parameterization
Previous work has usually assumed groundtruth access to relevant task-parameters [9, 8] or used infrared motion tracking devices [10], both of which restrict the applicability of the method. We therefore propose an approach to automatically select the relevant task parameters from RGB-D observations. Moreover, different skills in the same task require different information, e.g., in pick and place, the picking skill is often independent of the placing site. Modelling the full task using both task parameters is not only more computationally expensive but also leads to worse task performance. Hence, our parameter selection method has a synergistic effect with the skill segmentation proposed in Sec. IV-C. We offer a modular two-staged approach to candidate generation based on visual features and task-parameter selection based on statistical considerations. Both operate independently of each other, rendering them straightforward to substitute.
IV-F1 Candidate Generation
Various methods exist for extracting task parameters from visual observations [23]. Due to their generalization capabilities, we build on recent advances in 3D scene keypoints [4] and leverage a DINO vision transformer [12]. We sample a set of reference descriptors from the embedding of the first image observation of task demonstrations, either manually in a GUI or automatically using SAM-generated object masks [34]. The associated keypoints constitute the set of candidate task parameters. We then localize them in each demonstration by establishing correspondences via the embedding distances [4, 12].
IV-F2 Candidate Selection
We select task parameters per skill based on how well they explain the set of demonstrations. For each skill and candidate frame, we first fit a single-frame model. We only perform the time-based initialization described in Sec. IV-E and not the full EM to ensure temporal alignment across candidates and to speed up the procedure. We then estimate the relative relevance of the per-frame models using the determinants of their Gaussian precision matrices, a metric astutely proposed by Alizadeh et al. [11]. Due to the temporal alignment, the relative precision over Gaussian components yields the relative relevance of the candidate frames over time. By calculating the maximum over all Gaussian components, we select all frames that are relevant at some point in the segment. Let ${\scriptstyle\left(\boldsymbol{\Sigma}_{k}^{f}\right)^{-1}}$ denote the $k$-th component’s precision matrix in the model of frame $f$. For $C$ candidate frames with $K$ Gaussians, we estimate
$\omega(f)=\max_{k=1}^{K}\frac{\det\left(\boldsymbol{\Sigma}_{k}^{f}\right)^{-1}}{\sum_{c=1}^{C}\det\left(\boldsymbol{\Sigma}_{k}^{c}\right)^{-1}}$ (11)
and select all frames $f$ over some threshold $\omega(f)>\tau$.
In contrast, Alizadeh et al. fit a joint model to all candidates and the full task trajectory and calculate the relative precision during trajectory reconstruction [11]. Our per-skill single-frame approach scales to long-horizon tasks and large sets of candidate parameters while the temporal alignment eliminates the need for a reconstruction phase.
TABLE III: Time-driven vs. state-driven TAPAS-GMM, with and without disturbance of the rollout. Among the time-driven policies we compare different post-processing strategies Among the state-driven policies, we compare naive and factorized velocities. Trajectory length is mean and standard deviation.
Driver Variant Policy success rate $\uparrow$ Successful trajectory length [steps] $\downarrow$
Undisturbed Disturbed Undisturbed Disturbed
PushBlock LiftBottle PushBlock LiftBottle PushBlock LiftBottle PushBlock LiftBottle
Time None 0.85 0.18 0.00 0.00 $127\pm 01$ $239\pm 09$
Threshold 0.96 0.79 0.00 0.00 $149\pm 01$ $310\pm 06$
TOPP-RA 0.35 0.78 0.00 0.17 $370\pm 64$ $426\pm 26$ $444\pm 18$
State Naive 0.76 0.18 0.76 0.19 $206\pm 71$ $203\pm 36$ $266\pm 59$ $454\pm 37$
Factorized 0.83 0.33 0.83 0.32 $141\pm 53$ $223\pm 48$ $228\pm 47$ $471\pm 48$
TABLE IV: Real world policy success rates.
PickAndPlace PourDrink MakeCoffee SweepBlocks OpenCabinet StoreInDrawer BakeBread
LSTM 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Diffusion Policy 0.00 0.00 0.00 0.00 0.00 0.00 0.00
TP-GMM 0.20 0.08 0.00 0.16 0.00 0.68 0.36
TAPAS-GMM (Ours) 1.00 1.00 0.96 0.96 0.92 1.00 0.96
V Experimental Results
We first evaluate the efficacy of TAPAS-GMM for policy learning on a set of challenging manipulation tasks from the well established RLBench benchmark [13] and ablate the major components of our method. Afterward, we compare time-driven and state-driven policies on custom tasks implemented in ManiSkill2 [35]. Finally, we verify our results on a real robot and demonstrate our method’s generalization capabilities. Fig. 6 illustrates the evaluated RLBench tasks and real world tasks. Additional details for all the experiments are provided in Sec. S.4 of the supplementary material.
V-A Time-Driven Policy Learning in Simulation
We compare against LSTM [33] and Diffusion Policy [32] baselines trained with either 100 or 5 demonstrations. In contrast, TAPAS-GMM is only trained on five demonstrations. To experimentally disentangle policy learning and representation learning, we use groundtruth object poses as inputs for the baselines and as candidate task parameters for TAPAS-GMM. We evaluate our proposed visual features in Sec. V-C and Sec. V-D. We further compare against 3D Diffuser Actor [31], which at the time of writing is the incumbent on RLBench. Note that it leverages a 3D scene representation instead of object poses. We evaluate all policies for 200 episodes. Tab. I shows that TAPAS-GMM matches or outperforms the SOTA across all tasks while requiring significantly fewer demonstrations. Moreover, in contrast to all baselines, TAPAS-GMM scales to more complex long-horizon tasks such as SweepToDustpan and PlaceCups, without loss in performance. Moreover, unlike the Diffusion Policy and LSTM, it excels when high precision is required, e.g., in PushButton.
In Tab. II, we ablate the relevance of task-parameter selection and skill segmentation. The global TP selection chooses the parameters for the full task jointly, instead of per skill. As hypothesized, the task-parameter selection is more crucial on multi-object tasks with clutter (PlaceCups) than on single-object tasks (OpenDrawer, SlideBlock). Across all tasks, we observe that the initial end-effector frame is irrelevant after aligning the gripper with the object and reduces policy success if still taken into account. Skill segmentation is most crucial if the task is complex (PlaceCups) or precise grasps are required (OpenDrawer). The large performance gap between TAPAS-GMM and all other models on PlaceCups highlights the synergistic effect of skill segmentation and parameter selection.
V-B Time-Driven vs. State-Driven Policies
We compare the effectiveness of time-driven and state-driven GMMs in Tab. III. To simulate disturbances during policy rollout, we reset the end-effector pose back to the starting pose, around the object contact point in the trajectory. While state-based policies are harder to learn from few demonstrations, they outperform time-based policies when confronted with disturbances due to their time-invariance. Furthermore, our action-factorized policies consistently outperform the naive implementation.
Moreover, time-based GMMs do not take into account the velocity limits of the robot. Consequently, the predicted trajectories might not be feasible and require some post-processing, such as applying TOPP-RA [28] or only stepping the time when the current pose is close enough to the last prediction (thresholding). These post-processing methods have their own strengths and weaknesses. For example, TOPP-RA helps with precise grasping in LiftBottle and improves robustness against disturbances, but struggles with the drag of the object in PushBlock. In contrast, our state-driven GMMs explicitly model the robot’s velocity distribution in the demonstrations, thus not requiring post-processing. Often, they also produce shorter trajectories. The biggest drawback of state-driven GMMs that we observed experimentally, is their propensity for divergence during inference. Under velocity control, pose errors can accumulate and push the policy outside the observed state space. From this they often fail to recover, as frequently happens in LiftBottle.
V-C Real-World Policy Learning
We verify TAPAS-GMM capabilities extensively on a real Franka Emika robot. To this end, we construct a range of challenges such as long-horizon tasks up to 1000 time steps (e.g. MakeCoffee, BakeBread), tasks requiring high precision (e.g. MakeCoffee, BakeBread), and articulated objects (e.g. OpenCabinet, StoreInDrawer). All policies are trained on five task demonstrations and evaluated for 25 episodes. We generate candidate task parameters from wrist-camera observations as proposed in Sec. IV-F1.
The policy success rates in Tab. IV are consistent with the results in our simulated experiments and demonstrate the core advantages of our approach: 1. strong generalization from few demonstrations, 2. handling of articulated objects, and 3. scaling to complex tasks and long horizons. Consistent with our simulated ablation experiment, TP-GMMs struggle with pose accuracy and gripper timing due to locally irrelevant task parameters and poor temporal alignment. Moreover, the deep learning baselines fail to generalize from five demonstrations and do not learn to solve any task instance. In contrast, TAPAS-GMM effortlessly scales to long-horizon tasks, without any loss in accuracy or generalization performance.
V-D Visual Generalization and Skill Reuse
Prior work has extensively demonstrated the generalization capabilities of keypoints as visual features [4], which our policies inherit through the parameter generation proposed in Sec. IV-F1. We study the resulting zero-shot generalization capabilities of our policies on PickAndPlace and PourDrink by evaluating the policy with unseen objects instances, clutter objects, and task environments. We observe strong generalization performance across the board. We further provide qualitative results and discussions in the supplementary video.
Another advantage of TAPAS-GMM is that it not only solves one task but instead learns skill models, which can be recombined to solve novel tasks. In the supplementary video, we demonstrate how the segmented skills can be sequenced and reused in novel ways. In particular, the BakeBread task consists of taking the tray out of the oven and placing the bread on it. As an example, we invert and reorder these skills to construct a policy that takes the bread from the tray and inserts the tray into the oven.
VI Conclusion
We present TAPAS-GMM, a method for learning complex manipulation tasks from only five demonstrations. By segmenting skills and automatically selecting relevant task parameters per skill, our method scales both to tasks with long horizons and many parameters. Through leveraging visual semantic features, our method further generalizes across object instances, environments, and clutter. Finally, our proposed velocity factorization makes time-invariant task-parameterized policies more effective.
Limitations: TAPAS-GMM is vulnerable to some of the core problems of imitation learning. First, it does not reason about the robot’s kinematic limits or object collisions. Second, it requires a decent coverage of the task space to generalize, which can be difficult to achieve on multimodal trajectory distributions, when given only five demonstrations. This shows in the PlaceCups task, where workspace limits cause different task instances to require vastly different trajectory shapes. Adding a prior, encoding workspace limits, and object collision probabilities, might alleviate this problem.